OpenStudy (anonymous):

Select the inequality that corresponds to the given graph. 5x + 7 ≥2 and 10x ≤30 6x + 7 ≥ 25 or 2x - 8 < -4 4x + 10 ≤ 6 or 5x - 6 ≥ 14 37 ≤ 8x - 3 ≤ 61

OpenStudy (aravindg):

where is the graph ?

OpenStudy (anonymous):

@AravindG theres the graph

OpenStudy (whpalmer4):

Okay, we've got two filled-in circles, so our answer must have <= or >= in both inequalities. That eliminates the 2nd choice. Now look at the others and try some points. I like to use 0 as a test point, because the arithmetic is easy. Looking at the last choice, \(37 \le 8x -3 \le 61\) if we plug in 0, does it satisfy it? \(37 <= 8(0) -3) \le 61\) is false, so that can't be our answer. Repeat for the rest of the choices, until you find one that works.

OpenStudy (whpalmer4):

Nuts, that should be \[37 \le 8(0)-3 \le 61\]